# Resource-Efficient Hybrid Quantum-Classical Simulation Algorithm

### v1

@@ I Introduction
Quantum-Assisted Simulator (QAS) [40, 41] is a hybrid quantum-classical simulation algorithm that was originally proposed to combine NISQ and classical computers to simulate quantum dynamics of a system with a time-independent Hamiltonian. It assumes a time-evolution ansatz as a linear combination of basis states. Using a classical computer, the linear coefficients are evolved via a set of complex linear dynamical equations derived from dynamical variational principles. Although QAS utilizes dynamical variational principles for simulation, it does not have any quantum variational optimization subroutines [39]. This allows circumvention of the barren plateau problem [42, 43] as there is no need for any gradient measurements [40, 41]. Moreover, it guarantees energy conservation for simulations with time-independent Hamiltonians, that are distinct from most other variational quantum simulators in the literature [44, 45, 46, 47, 48, 49, 50]. The quantum computer is only used to calculate basis state overlaps, the Hamiltonian elements as defined in the dynamical equations and also the observable elements expressed in the basis states. The dynamical properties of the quantum state are then classically computed using the observable elements and the time-evolved linear coefficients. This avoids the computational bottleneck of preparing multiple copies of time-evolved quantum states for measurement at every intermediate time step.
@@
In this work, we shall show how QAS can leverage FTQC devices and quantum time propagation algorithms to simulate quantum dynamics efficiently by considering time-evolved states themselves as a basis, as shown in Fig 1b. The time-evolved basis set size thus becomes equal to the time-evolution Hilbert subspace, where all time-evolved states reside. Specifically, if the system is initialized in an unknown superposition of $n$ system eigenstates, the QAS will only need to store and evolve $n$ linear coefficients, without needing any information on the eigenvalues and eigenstates. For a $2N$-qubit quantum device, QAS is expected to consume fewer quantum resources than the standard method when the number of time steps exceeds roughly $100n^{2}$. Therefore, if $n{\ll}2^{N}$ is small, then QAS is able to perform more efficiently in simulating large $N$ system sizes, for a larger number of time steps than most existing hybrid quantum simulation algorithms. As a simple example, we shall demonstrate QAS for simulating quantum dynamics for a 4-qubit Helium atom and 8-qubit Hydrogen molecule initialized in a superposition of $n{=}2$ eigenstates.
Figure 1: (a) Standard method to simulate the quantum dynamics of a system of Hamiltonian $\hat{H}$, initialized in $|\psi_{0}\rangle$, up to a time $T$, measured by an observable $\hat{O}$ at fixed time intervals $\Delta t$. (b) Our proposal to use quantum-assisted simulator to assume a time evolution state $|\psi(t)\rangle$ as a linear combination of time-evolved states. We assumed atomic units, $\hbar{=}1$, in this figure.
@@ II.1 Quantum-Assisted Simulator
QAS is a hybrid quantum-classical simulation algorithm that considers a time-evolution ansatz |$\ket{\psi(t)}$ as a complex linear combination of basis states {|$\{\ket{\psi_{j}}\}$ that is,
|$\ket{\psi(t)}=\sum_{j=0}\alpha_{j}(t)\ket{\psi_{j}},$ (2)
@@
$\delta_{\boldsymbol{\alpha}}\|(i\hbar\partial_{t}-\hat{H})\ket{\psi(t)}\|=0$ (3)
@@
$\boldsymbol{F}\dot{\boldsymbol{\alpha}}=-i\boldsymbol{H}\boldsymbol{\alpha}$ (4)
@@
    $\braket{\psi(t)}{\hat{O}}{\psi(t)}=\sum_{jk}\alpha^{*}_{j}(t)O_{jk}\alpha_{k}(t).$ (7)
@@ II.2 Minimizing Quantum-Classical Resources
The hybrid quantum-classical setup in QAS enables certain types of simulations to run efficiently. In particular, consider a system of size $N$ initialized in the following unknown superposition of $n$ non-degenerate eigenstates of the system |$\ket{e_{j}}$, where $n\ll 2^{N}$.
@@
where $s_{j}$ are parameter times which are not more than the total simulation time, that is $0{=}s_{0}{<}s_{1}{<}{\ldots}{<}s_{n-1}{\leq}T$, using atomic units $\hbar{=}1$, which shall be used throughout this paper. Then, assuming linear independence of the basis, we will only need $n$ time-evolved basis states that fully span the time-evolution Hilbert subspace. The quantum computer then handles these basis states with a dimension of $2^{N}$ each, whilst the classical computer will only need to store and evolve just $n$ linear coefficients. The system eigenvalues, eigenstates |$\ket{e_{j}}$, and the linear coefficients $\beta_{j}$ in the initial state in Eq. (8) remain unknown to us and need not be solved.
Assuming the system Hamiltonian $\hat{H}$ and observable $\hat{O}$ both decompose into a linear combination of $L$ number of Pauli strings $\hat{P}_{l}{=}\otimes^{N}_{j=1}\hat{\sigma}_{j}$ where $\hat{\sigma}_{j}{\in}\{\hat{I}_{j},\hat{X}_{j},\hat{Y}_{j},\hat{Z}_{j}\}$, then getting the basis state overlap $\boldsymbol{F}$, Hamiltonian $\boldsymbol{H}$ and observable $\boldsymbol{O}$ matrices requires the estimation of the following quantities
@@
$P_{jkl}$ $=\braket{\psi_{0}}{\hat{P}_{l}e^{i\hat{H}\Delta s_{jk}}}{\psi_{0}}$ (11)
where $\Delta s_{jk}{=}s_{j}{-}s_{k}$ parameterises the differences in the basis states’ evolution times.
We first consider using the Hadamard Test with the time propagator $U{=}e^{-i\hat{H}\Delta t}$ to evaluate $F_{jk}$ and $P_{jkl}$. An ancilla-controlled-$U$ would have longer quantum runtime complexity than the standard $U$, but by no more than a constant multiple factor $\gamma$ [56]. If we now assume access to a $2N$-qubit quantum computer, we can consider a modified Hadamard test with $N$ ancilla qubits [56], instead of just one ancilla qubit considered previously. We refer readers to Appendix A for more details on the modified Hadamard Test. Controlled-quantum gates in the controlled-$U$ that act on separate sets of systems qubits can remain parallel if the gates are controlled by separate ancilla qubits. In the worst case scenario, since all times $s_{j}$ are not more than the total simulation time $T$, all modified Hadamard tests have a maximum number of $\frac{T}{\Delta t}$ ancilla controlled-$U$s. The total overall quantum runtime complexity to estimate all the required quantities in Eqs. (10) and (11) is thus $\mathcal{O}\left(\frac{T\gamma n^{2}}{\Delta t}\textrm{Poly}(L,\epsilon^{-1},\Delta t)\right)$ where $\textrm{Poly}(L,\epsilon^{-1},\Delta t)$ is the quantum runtime complexity of $U$ and $\epsilon{\geq}\left\|e^{i\hat{H}\Delta t}{-}\hat{U}\right\|$ is the time evolution error.
In contrast, the standard method of preparing multiple copies of the time-evolved states at every intermediate time step requires $\mathcal{O}\left(\left(\frac{T}{\Delta t}\right)^{2}\right)$ repetitions of $U$. The overall quantum runtime complexity to run the standard method is thus $\mathcal{O}\left(\left(\frac{T}{\Delta t}\right)^{2}\textrm{Poly}(L,\epsilon^{-1},\Delta t)\right)$. We therefore estimate that QAS will consume fewer quantum resources than the standard method in terms of overall quantum runtime complexity, when the number of time steps exceeds $\frac{T}{\Delta t}{\gtrsim}16\gamma n^{2}$. By making a reasonable assumption of $\gamma{\approx}6$, based on the general observation that a 3-qubit Toffoli gate can be decomposed into at least 6 CNOT gates [57, 58], the threshold is approximated to be $\frac{T}{\Delta t}{\gtrsim}100n^{2}$. We refer readers to the full derivation of this claim in Appendix B.
@@ III Examples and Results
$\hat{H}_{\mathrm{elec}}=\sum_{pq}^{N}h_{pq}\hat{a}^{\dagger}_{p}\hat{a}_{q}+\frac{1}{2}\sum_{pqrs}^{N}h_{pqrs}\hat{a}^{\dagger}_{p}\hat{a}^{\dagger}_{q}\hat{a}_{r}\hat{a}_{s}$ (12)
@@
We implemented a numerical statevector calculation using the Jordan-Wigner (JW) fermion-to-qubit mapping [60], which transforms $\hat{H}_{\mathrm{elec}}$ into a Pauli Hamiltonian $\hat{H}_{P}$ with up to $L{\sim}O(N^{4})$ terms in general [61, 4]. $\hat{H}_{\mathrm{elec}}$ for the He atom is mapped to a 4-qubit system with 27 terms in $\hat{H}_{P}$, and that for H2 molecule is mapped to an 8-qubit system with 185 terms in $\hat{H}_{P}$. The time-evolved basis set consist of $n{=}2$ states: the initial state |$\ket{\psi_{0}}$ and |$\ket{\psi_{1}}{=}e^{-i\hat{H}/2}\ket{\psi_{0}}$, where we set the parameter time $s_{1}{=}\tfrac{1}{2}$. Assuming an ideal noiseless Hadamard Test with $10^{4}$ simulated shots, we randomly sampled 100 sets of the basis state overlaps $\boldsymbol{F}$, Hamiltonian matrices $\boldsymbol{H}$ and the orbital population observable $\boldsymbol{O}$. Next, we solve the ODE according to Eq. (4) for every sample pair of $\boldsymbol{F}$ and $\boldsymbol{H}$ independently up to a simulation time $t{=}4\textrm{ Hartree}^{-1}$ with a time interval of $\Delta t{=}0.001\textrm{ Hartree}^{-1}$. The total number of time steps in this simulation is $4000$, which far exceeds the threshold of $400100\cdot 2^{2}{=}$ 400 time steps, placing this QAS demonstration well into quantum resource-efficient regime. Finally, we calculated the orbital population at every time step for each sample run. We refer readers to Appendix C for details on how the sampling of $\boldsymbol{F}$, $\boldsymbol{H}$ and $\boldsymbol{O}$ matrices was performed.
@@
We plotted the orbital population dynamics of the He atom and H2 molecule in Figs. 2 and 3 respectively. The colored solid lines and shaded regions represents the population mean and uncertainty of the 100 independent QAS simulation samples, respectively. The true time-evolution is denoted by a black dashed line. In addition, we also refer readers to Appendix D for plots of the total, coulomb, kinetic and potential energy dynamics, which demonstrate the total energy conservation feature of QAS. For both systems, we observe the orbital population oscillates with a frequency of roughly 0.5 Hartree. This is in agreement with the true frequency of -0.554 and -0.490 Hartrees, respectively. For the helium atom, the maximum fractional population uncertainty is about 4% for the $1s$ and 5% for the $2s$ atomic orbitals. For the hydrogen molecule, the maximum fractional population uncertainty is about 6% for both $1\sigma$ and $2\sigma*$ molecular orbitals. We also observe that the population estimation of $1\sigma*$ and $2\sigma$ molecular orbitals are extremely imprecise due to its low orbital population. The low precision can be improved by simply increasing number of shots, for which we refer reader to Appendix E which details the linear relationship between the variance of the estimated quantities and number of shots. In principle, if the quantum subroutines are executed with minimal and unbiased quantum noise, that is indistinguishable to statistical shot noise, then the QAS can accurately simulate the quantum dynamics, though more shots are needed to improve its precision.
@@ IV Discussion and Outlook
Our results have demonstrated that useful dynamical quantities of chemical systems can be estimated fairly accurately and precisely with modest amount of quantum-classical resources and minimal heuristics. However, the QAS performance depends heavily on the quality of the quantum time propagation algorithm employed. As an example, we refer readers to Appendix F where we show how a large number of Trotter steps from the first-order Trotterization of time-evolved basis is required to achieve good simulation fidelity for the helium atom case. The trade-offs between the quality of the time-evolved basis state and the overall quantum runtime complexity must be thoroughly analyzed. Thus, an important direction would be to identify and refine time propagation techniques that would minimize quantum resources while maximizing simulation accuracy for different system types.
Whilst QAS may seem very accurate when compared with other simulation algorithms, there is a major caveat that needs to be pointed out. The time-evolved basis is often linearly dependent, making the complex dynamical equation in Eq. (4) prone to ill-conditioning, thus resulting in numerical instabilities. As the number of eigenstates in the initial state $n$ increases, there is an increasing likelihood of the time-evolved basis becoming too similar to each other as the corresponding eigenbasis amplitudes only differ in its phase angles but not in its magnitude, resulting in linear dependence of the basis. The parameter times must then be chosen carefully to avoid such an issue. In the $n{=}2$ base case, any parameter time $s_{1}{\neq}\tfrac{2\pi k}{\Delta e}$, for any $k{\in}\mathbb{Z}$ and eigenvalue difference $\Delta e$, is acceptable as it leads to a well-conditioned problem, as shown in Appendix G. For the $n{>}2$ case, however, it becomes an incredibly difficult problem and remains an open question on the best method to choose a optimal set of parameter times. In practice, for small $n$, parameter times are chosen heuristically via trial and error to ensure to an accurate simulation.

### v2

@@ I Introduction
Quantum-Assisted Simulator (QAS) [40, 41] is a hybrid quantum-classical simulation algorithm that was originally proposed to combine NISQ and classical computers to simulate quantum dynamics of a system with a time-independent Hamiltonian. It assumes a time-evolution ansatz as a linear combination of basis states. Using a classical computer, the linear coefficients are evolved via a set of complex linear dynamical equations derived from dynamical variational principles. Although QAS utilizes dynamical variational principles for simulation, it does not have any quantum variational optimization subroutines [39]. This allows the circumvention of the barren plateau problem [42, 43] as there is no need for any estimation of the variational landscape [40, 41]. Moreover, it guarantees energy conservation for simulations with time-independent Hamiltonians, that are distinct from most other variational quantum simulators in the literature [44, 45, 46, 47, 48, 49, 50]. The quantum computer is only used to calculate basis state overlaps, the Hamiltonian elements as defined in the dynamical equations and also the observable elements expressed in the basis states. The dynamical properties of the quantum state are then classically computed using the observable elements and the time-evolved linear coefficients. This avoids the computational bottleneck of preparing multiple copies of time-evolved quantum states for measurement at every intermediate time step.
@@
In this work, we shall show how QAS can complement and utilize quantum time propagation algorithms implemented on FTQC devices to simulate quantum dynamics efficiently by considering time-evolved states themselves as a basis, as shown in Fig 1b. The time-evolved basis set size thus becomes equal to the time-evolution Hilbert subspace, where all time-evolved states reside. Specifically, if the system is initialized in an unknown superposition of $n$ system eigenstates, the QAS will only need to store and evolve $n$ linear coefficients, without needing any information on the eigenvalues and eigenstates. For a $2N$-qubit quantum device, QAS is expected to consume fewer quantum resources than the standard method when the number of time steps exceeds roughly $100n^{2}$. Therefore, if $n{\ll}2^{N}$ is small, then QAS is able to perform more efficiently in simulating large $N$ system sizes, for a larger number of time steps than most existing hybrid quantum simulation algorithms. As a simple example, we shall demonstrate QAS for simulating quantum dynamics for a 4-qubit Helium atom and 8-qubit Hydrogen molecule initialized in a superposition of $n{=}2$ eigenstates.
Figure 1: (a) Standard method to simulate the quantum dynamics of a system of Hamiltonian $\hat{H}$, initialized in $|\psi_{0}\rangle$, up to a time $T$, measured by an observable $\hat{O}$ at fixed time intervals $\Delta t$. (b) Our proposal to use Quantum-Assisted Simulator (QAS) to utilize quantum time propagation algorithms to efficiently simulate quantum dynamics. QAS assumes a time evolution state $|\psi(t)\rangle$ as a linear combination of time-evolved states. Atomic units, $\hbar{=}1$, are assumed in this figure.
@@ II.1 Quantum-Assisted Simulator
QAS is a hybrid quantum-classical simulation algorithm that considers a time-evolution ansatz |$\ket{\psi(t)}$ as a complex linear combination of $n_{\textrm{basis}}$ basis states {|$\{\ket{\psi_{j}}|j{=}0,1,{\ldots},n_{\textrm{basis}}{-}1\}$ that is,
|$\ket{\psi(t)}=\sum_{j=0}^{n_{\textrm{basis}}-1}\alpha_{j}(t)\ket{\psi_{j}},$ (2)
@@
$\delta_{\boldsymbol{\alpha}}\|(i\hbar\partial_{t}-\hat{H})\ket{\psi(t)}\|=0,$ (3)
@@
$\boldsymbol{F}\dot{\boldsymbol{\alpha}}=-i\boldsymbol{H}\boldsymbol{\alpha},$ (4)
@@
We refer readers to Appendix A for detailed derivation of parameter evolution in Eq. (4).
@@
    $\braket{\psi(t)}{\hat{O}}{\psi(t)}=\sum_{j,k=0}^{n_{\textrm{basis}}-1}\alpha^{*}_{j}(t)O_{jk}\alpha_{k}(t).$ (7)
@@ II.2 Minimizing Quantum-Classical Resources
The hybrid quantum-classical setup in QAS enables certain types of simulations to run efficiently. In particular, consider a system of size $N$ initialized in the following unknown superposition of $n$ non-degenerate eigenstates of the system |$\ket{e_{j}}$, where $n$ is defined as the number of eigenstates that span the time-evolution subspace of the quantum system, and assuming $n\ll 2^{N}$,
@@
where $s_{j}$ are parameter times which are not more than the total simulation time, that is $0{=}s_{0}{<}s_{1}{<}{\ldots}{<}s_{n_{\textrm{basis}}-1}{\leq}T$, using atomic units $\hbar{=}1$, which shall be used throughout this paper. Then, assuming linear independence of the basis, we will only need $n_{\textrm{basis}}{=}n$ time-evolved basis states that fully span the time-evolution Hilbert subspace. The quantum computer then handles these basis states with a dimension of $2^{N}$ each, whilst the classical computer will only need to store and evolve just $n_{\textrm{basis}}{=}n$ linear coefficients. If there are fewer time-evolved basis states than $n$, that is $n_{\textrm{basis}}{<}n$, the time-evolution ansatz is under-parameterized, leading to incorrect results. On the other hand, if the number of basis states is much larger than $n$, that is $n_{\textrm{basis}}{>}n$, the ansatz becomes over-parameterized which may lead to numerical convergence issues. The system eigenvalues, eigenstates |$\ket{e_{j}}$, and the linear coefficients $\beta_{j}$ in the initial state in Eq. (8) remain unknown to us and need not be solved.
Assuming the system Hamiltonian $\hat{H}$ and observable $\hat{O}$ both decompose into a linear combination of $L$ number of Pauli strings $\hat{P}_{l}{=}\otimes^{N}_{j=1}\hat{\sigma}_{j}$ where $\hat{\sigma}_{j}{\in}\{\hat{I}_{j},\hat{X}_{j},\hat{Y}_{j},\hat{Z}_{j}\}$, which may or may not share the same Pauli string, then getting the basis state overlap $\boldsymbol{F}$, Hamiltonian $\boldsymbol{H}$ and observable $\boldsymbol{O}$ matrices requires the estimation of the following quantities
@@
$P_{jkl}$ $=\braket{\psi_{0}}{\hat{P}_{l}e^{i\hat{H}\Delta s_{jk}}}{\psi_{0}},$ (11)
where $\Delta s_{jk}{=}s_{j}{-}s_{k}$ parameterizes the differences in the basis states’ evolution times.
We first consider using the Hadamard Test with the time propagator $\hat{U}{=}e^{-i\hat{H}\Delta t}$ to evaluate a combined total of $\mathcal{O}(Ln^{2})$ quantities, $F_{jk}$ and $P_{jkl}$ in Eqs. (10) and (11) respectively. An ancilla-controlled-$\hat{U}$ would have longer a quantum runtime complexity than the standard $\hat{U}$, but by no more than a constant multiple factor $\gamma$ [56]. If we now assume access to a $2N$-qubit quantum computer, we can consider a modified Hadamard test with $N$ ancilla qubits [56], instead of just one ancilla qubit considered previously. We refer readers to Appendix B for more details on the modified Hadamard Test. Controlled-quantum gates in the controlled-$\hat{U}$ that act on separate sets of systems qubits can remain parallel if the gates are controlled by separate ancilla qubits. In the worst case scenario, since all times $s_{j}$ are not more than the total simulation time $T$, all modified Hadamard tests have a maximum number of $\frac{T}{\Delta t}$ ancilla controlled-$U$s. The total overall quantum runtime complexity to estimate all the required quantities is thus $\mathcal{O}\left(\frac{T\gamma Ln^{2}}{\Delta t}\textrm{Poly}(L,\epsilon^{-1},\Delta t)\right)$ where $\textrm{Poly}(L,\epsilon^{-1},\Delta t)$ is the quantum runtime complexity of $\hat{U}$ and $\epsilon{\geq}\left\|e^{i\hat{H}\Delta t}{-}\hat{U}\right\|$ is the time evolution error.
In contrast, the standard method involves preparing $L$ copies of the time-evolved states and measuring each with the corresponding Pauli strings, decomposed from the observable, at every intermediate time step. Thus, this requires $\mathcal{O}\left(L\left(\frac{T}{\Delta t}\right)^{2}\right)$ repetitions of $\hat{U}$. The overall quantum runtime complexity to run the standard method is thus $\mathcal{O}\left(L\left(\frac{T}{\Delta t}\right)^{2}\textrm{Poly}(L,\epsilon^{-1},\Delta t)\right)$. We therefore estimate that QAS will consume fewer quantum resources than the standard method in terms of overall quantum runtime complexity, when the number of time steps exceeds $\frac{T}{\Delta t}{\gtrsim}16\gamma n^{2}$. By making a reasonable assumption of $\gamma{\approx}6$, based on the general observation that a 3-qubit Toffoli gate can be decomposed into at least 6 CNOT gates [57, 58], the threshold is approximated to be $\frac{T}{\Delta t}{\gtrsim}100n^{2}$. We refer readers to the full derivation of this claim in Appendix C.
@@ III Examples and Results
$\hat{H}_{\mathrm{elec}}=\sum_{pq}^{N}h_{pq}\hat{a}^{\dagger}_{p}\hat{a}_{q}+\frac{1}{2}\sum_{pqrs}^{N}h_{pqrs}\hat{a}^{\dagger}_{p}\hat{a}^{\dagger}_{q}\hat{a}_{r}\hat{a}_{s},$ (12)
@@
We implemented a numerical statevector calculation using the Jordan-Wigner (JW) fermion-to-qubit mapping [60], which transforms $\hat{H}_{\mathrm{elec}}$ into a Pauli Hamiltonian $\hat{H}_{P}$ with up to $L{\sim}O(N^{4})$ terms in general [61, 4]. $\hat{H}_{\mathrm{elec}}$ for the He atom is mapped to a 4-qubit system with 27 terms in $\hat{H}_{P}$, and that for H2 molecule is mapped to an 8-qubit system with 185 terms in $\hat{H}_{P}$. The time-evolved basis set consist of $n{=}2$ states: the initial state |$\ket{\psi_{0}}$ and |$\ket{\psi_{1}}{=}e^{-i\hat{H}/2}\ket{\psi_{0}}$, where we set the parameter time $s_{1}{=}\tfrac{1}{2}$. Assuming an ideal noiseless Hadamard Test with $10^{4}$ simulated shots, we randomly sampled 100 sets of the basis state overlaps $\boldsymbol{F}$, Hamiltonian matrices $\boldsymbol{H}$ and the orbital population observable $\boldsymbol{O}$. Next, we solve the ODE according to Eq. (4) for every sample pair of $\boldsymbol{F}$ and $\boldsymbol{H}$ independently up to a simulation time $t{=}4\textrm{ Hartree}^{-1}$ with a time interval of $\Delta t{=}0.001\textrm{ Hartree}^{-1}$. The total number of time steps in this simulation is $4000$, which far exceeds the threshold of $400100\cdot 2^{2}{=}$ 400 time steps, placing this QAS demonstration well into quantum resource-efficient regime. Finally, we calculated the orbital population at every time step for each sample run. We refer readers to Appendix D for details on how the sampling of $\boldsymbol{F}$, $\boldsymbol{H}$ and $\boldsymbol{O}$ matrices was performed.
@@
We plotted the orbital population dynamics of the He atom and H2 molecule in Figs. 2 and 3 respectively. The colored solid lines and shaded regions represents the population mean and uncertainty of the 100 independent QAS simulation samples, respectively. The true time-evolution is denoted by a black dashed line. In addition, we also refer readers to Appendix E for plots of the total, coulomb, kinetic and potential energy dynamics, which demonstrate the total energy conservation feature of QAS. For both systems, we observe the orbital population oscillates with a frequency of roughly 0.5 Hartree. This is in agreement with the true frequency of -0.554 and -0.490 Hartrees, respectively. For the helium atom, the maximum fractional population uncertainty is about 4% for the $1s$ and 5% for the $2s$ atomic orbitals. For the hydrogen molecule, the maximum fractional population uncertainty is about 6% for both $1\sigma$ and $2\sigma*$ molecular orbitals. We also observe that the population estimation of $1\sigma*$ and $2\sigma$ molecular orbitals are extremely imprecise due to its low orbital population. The low precision can be improved by simply increasing number of shots, for which we refer reader to Appendix F which details the linear relationship between the variance of the estimated quantities and number of shots. In principle, if the quantum subroutines are executed with minimal and unbiased quantum noise, that is indistinguishable to statistical shot noise, then the QAS can accurately simulate the quantum dynamics, though more shots are needed to improve its precision.
@@ IV Discussion and Outlook
Our results have demonstrated that useful dynamical quantities of chemical systems can be estimated fairly accurately and precisely with modest amount of quantum-classical resources and minimal heuristics. However, the QAS performance depends heavily on the quality of the quantum time propagation algorithm employed. As an example, we refer readers to Appendix G where we show how a large number of Trotter steps from the first-order Trotterization of time-evolved basis is required to achieve good simulation fidelity for the helium atom case. The trade-offs between the quality of the time-evolved basis state and the overall quantum runtime complexity must be thoroughly analyzed. Thus, an important direction would be to identify and refine time propagation techniques that would minimize quantum resources while maximizing simulation accuracy for different system types.
Whilst QAS may seem very accurate when compared with other simulation algorithms, there is a major caveat that needs to be pointed out. The time-evolved basis is often linearly dependent, making the complex dynamical equation in Eq. (4) prone to ill-conditioning, thus resulting in numerical instabilities. As the number of eigenstates in the initial state $n$ increases, there is an increasing likelihood of the time-evolved basis becoming too similar to each other as the corresponding eigenbasis amplitudes only differ in its phase angles but not in its magnitude, resulting in linear dependence of the basis. The parameter times must then be chosen carefully to avoid such an issue. In the $n{=}2$ base case, any parameter time $s_{1}{\neq}\tfrac{2\pi k}{\Delta e}$, for any $k{\in}\mathbb{Z}$ and eigenvalue difference $\Delta e$, is acceptable as it leads to a well-conditioned problem, as shown in Appendix H. For the $n{>}2$ case, however, solving for the complete set of conditions for the parameter times becomes an incredibly difficult problem and remains an open question on the best method to choose a optimal set of parameter times. In practice, for small $n$, a suitable set of parameter times is chosen heuristically via trial and error to avoid the complexity of solving the conditions so as to ensure an accurate and efficient simulation.